### v1

# Predicting Field Experiments with Large Language Models

Abstract
Large language models (LLMs) have demonstrated unprecedented emergent capabilities, including content generation, translation, and the simulation of human behavior. Field experiments, despite their high cost, are widely employed in economics and the social sciences to study real-world human behavior through carefully designed manipulations and treatments. However, whether and how these models can be utilized to predict outcomes of field experiments remains unclear. In this paper, we propose and evaluate an automated LLM-based framework that produces predictions of field experiment outcomes. Applying this framework to 319 experiments drawn from renowned economics literature yields a notable prediction accuracy of 78%. Interestingly, we find that performance is highly skewed. We attribute this skewness to several factors, including gender differences, ethnicity, and social norms.
1 Introduction
Field experiments allow researchers to manipulate variables of interest in a real-world setting, establishing causal relationships between interventions and outcomes. They typically begin by designing an intervention aligned with a specific research question, randomly assigning participants to treatment or control groups, and then measuring outcomes under natural conditions. The resulting data are collected and analyzed to evaluate the causal impact of the intervention Levitt and List (2009). It is adopted by a wide range of disciplines across academia and industry such as human behavior understanding Kao et al. (2024), marketing strategies making Hsu and Karahalios (2024), impact of machine learning on human Reinmund et al. (2024).
With the advancement of Generative AI, many scholars believe it is possible to leverage LLMs’ content generation and reasoning ability to simulate human responses and behavior Taubenfeld et al. (2024). Existing papers have confirmed the alignments between simulated data generated by LLMs and real data collected from human participants across various aspects, including human responses, traits Dubois et al. (2024), moral standards Dillion et al. (2023), preferences Namikoshi et al. (2024), and emotions Gao et al. (2023). Based on those encouraging results, scholars have successfully steered LLMs to replicate existing lab experiments across several disciplines, including psychology Aher et al. (2023), sociology Leng and Yuan (2023); Manning et al. (2024), and economics Aher et al. (2023); Horton (2023), which aim to replicate existing lab experiments by treating LLMs as participants of lab experiments.
However, these works have several disadvantages. First, these works only focused on lab experiment settings. But field experiments are inherently more challenging to conduct than laboratory experiments, due to diverse participant backgrounds, complex workflows, and multifaceted treatment designs. Second, previous studies mainly replied on manual process and only showcased a handleful experiments most related to social psychology. In order to examine LLMs’ capabilities robustness and generalizability in experiment simulation, we need to test it against a much larger scale and broader range of experiments.
In this paper, we fulfill the literature gap by proposing an automated LLM framework based on GPT to directly predict the outcome of field experiments. This framework has several components: an information extraction module, a variant generation module, and a prediction module. We test this framework on over 300 field experiments containing a total of 1638 conclusions. Using an LLM-powered multilevel filtering mechanism, we get these experiments screening from over 6,500 papers related to “field experiment” published in premier academic journals from 2000 to 2024.
Without any alignment techniques or fine-tuning, our framework achieves an average prediction accuracy of 78%. Specifically, for each conclusion, the framework generates two false variants, then inserts the true conclusion and two distractors into a prediction prompt template with a Chain-of-Thought design, prompting the LLM to select one of the three options. Interestingly, we also find that the prediction results are highly skewed. For example, our framework achieves nearly 100% prediction accuracy on 73% of conclusions while it completely fails to predict 16% of the conclusions with close to 0% accuracy. Further investigation on that skewness reveals that the LLM-based framework has limitations on predicting experiment with a context of gender differences, ethnicity, and social norms.
Our research makes several contributions. First, we extend the literature on LLMs’ emergent capabilities by demonstrating that LLMs can simulate field experiments by predicting conclusions. While using LLMs to simulate human behavior is well studied, to the best of our knowledge, this is the first work in literature that considers simulating field experiments that require a more complex environment setting and workflow design.
Second, our proposed framework allows the prediction of field experimental conclusions in a fully automated, large-scale fashion. As a result, our framework’s capability is more convincing and can be generalized to a wider range of applications.
Third, by comparing the predictions and ground truth from existing field experiments, this paper identifies infeasible scenarios where the LLM underperforms in prediction tasks, revealing the limitations of the LLM in field experimental simulations caused by confound reasons.
2 Literature Review
Our study is related to the LLM literature on simulating human behavior, with a particular focus on experimental simulations by LLMs. Here, we highlight our contributions by comparing and contrasting our work with existing studies.
While past literature mainly focused on agent-based social simulation Leung et al. (2024); Tennant et al. (2023); Zehmakan et al. (2023), there is an increasing trend to adopt LLMs as simulation tools. Existing studies have found that LLMs’ ability to simulate human behavior source from their possession of human-like reasoning skills and their adaptivity to personas of diverse characters Shanahan (2024); Taubenfeld et al. (2024). Upon those features of LLMs, Aher et al. Aher et al. (2023) proposed a concept of Turing Experiment that LLMs can be profiled as synthetic participants of experiments with integrated prompt of experimental settings and demographic information. Similarly, Horton Horton (2023) proved LLMs’ ability to simulate lab experiments and promoted it as a method of experimental pilot testing. The usage of LLMs in Horton Horton (2023)’s work is similar to Aher et al. Aher et al. (2023), consisting of two stages: prompting LLMs as synthetic experimental participants and collecting responses from conversations, reporting a few successful simulation cases of lab experiments. Leng & Yuan Leng and Yuan (2023) harnessed a three-phase procedure to complete the lab experimental simulation, which includes the initialization phase, interaction phase, and decision analysis phase. In the initialization phase, separate conversations of GPT-4 are prompted as vanilla experimental participants without specifying demographic information. Then, in the interaction stage, synthetic participants are prompted with the actions of other participants and asked to take actions and rationales according to the experiment design. Last, agents’ actions and rationales are collected for analysis to conclude. Leveraging this procedure, the study simulates five existing lab experiments. Manning et al. Manning et al. (2024) proposed an automated framework for various lab experimental simulations. Although the entire workflow is divided into seven steps, it is essentially similar to the manual procedures of the three papers above. However, the sole input is the backstory of the experiment. Based on the input information and powered by an LLM, the framework continues with subsequential steps: identifying dependent and independent variables, generating treatment values, profiling agents, organizing the interaction of agents, collecting data, and establishing causal relationships and conclusions. The paper reports the results of four social scenarios.
In summary, while existing studies discussed above pioneered the application of Large Language Models (LLMs) to laboratory experiment simulations, several research gaps remain unaddressed. For instance, existing methods are not tested on field experiments that are inherently more complex than laboratory experiments, encompassing diverse participant backgrounds, more intricate workflows, and multifaceted treatment designs. Besides, none of these studies examined the impact of bias on simulation outcomes, despite extensive discussions of AI bias at IJCAI Wang et al. (2023); Du and Zhang (2024). Given that recognized biases in LLMs—such as gender Khorramrouz et al. (2023) and social norm bias Ren et al. (2024)—can compromise performance on downstream tasks Harris et al. (2022), it is crucial to identify scenarios in which bias undermines the fidelity of experimental simulations. Another limitation is that these studies tested their simulation strategies on only a small number of experiments, restricting generalizability. To address these gaps, we evaluated our proposed framework at scale on field experiments from published papers, demonstrating not only that LLMs can accurately predict the outcomes of established experiments, but also clarifying boundary conditions under which they cannot provide reliable experimental predictions. Furthermore, the only existing automated simulation framework Manning et al. (2024) is incompatible with scenarios involving complex human-object interactions or precise experimental control. By contrast, our framework offers a broadly adaptable approach that supports experiments across a wide range of contexts.
3 Data Collection and Filtering
To test LLMs’ ability to predict field experiments in a large-scale fashion, we collected existing field experiments from papers published in the top 4 journals in economics, such as The Review of Economic Studies, American Economic Review, Journal of Political Economy, and The Quarterly Journal of Economics. It is worth noting that we focus on field experiments in economics because they are generally with a larger scale and reflect more generic human behaviors compared to field experiments in other disciplines Levitt and List (2009).
Figure 1: The Data Collection Workflow
Figure 1 shows the data collection and filtering workflow. Initially, 6544 papers containing keywords related to field experiments published between 2000 and 2024 were selected. Then, we applied a two-layer verification process powered by Claude (Claude-3-opus-20240229). First, we prompted the title and abstract of each paper to Claude and asked it to judge if the paper designs and implements a field experiment. Upon this, the second verification prompted the entire paper to Claude, asking the same question. This strategy balances the accuracy of verification and the cost of calling Claude, as prompting the entire paper exponentially increases the cost. It is also worth mentioning that using Claude as the verification tool, instead of GPT, prevents potential data leakage as GPT is leveraged as the prediction tool in our framework. The distribution of the selected papers is shown in Figure 2.
Figure 2: A Summary of Qualified Papers
4 Framework
We present the details of our automated framework for predicting field experimental conclusions in this section. Figure 3 shows the workflow. Overall, our framework is divided into three stages: Extraction from Papers, Variant Generation, and Prediction. Notably, Claude (Claude-3-opus-20240229) powers all preprocess tasks in the first two stages, whereas GPT completes the prediction at the last stage. We use two different LLMs in different stages to prevent potential data leakage from one another.
Figure 3: Prediction Framework
4.1 Extraction
Figure 4: Prompt for Extraction
Specicifically, the framework uses Claude to extract information related to a field experiment from the selected paper. To realize this, the framework leverages a manually crafted prefixed prompt, which has proven to be efficient for various downstream tasks Brown et al. (2020). As shown in Figure 4, the prompt template contains a placeholder “Paper”, an information form consisting of bullet points from “A” to “G”, and clear instructions that ask the LLM to extract information according to the form from the paper. As underscored in the prompt, the first six bullet points “A” to “F” are key experimental settings that shape the experiment context, whereas the last point “G” is about conclusions that are true outcomes in the prediction task.
Based on the response from Claude, the framework formulates “Experiment Settings” directly from bullet points “A” to “F”, while it polishes point “G” to generate “Experiment Conclusions”. Specifically, raw response regarding point “G” is a paragraph containing multiple conclusions of the field experiment. To separate that paragraph into standalone conclusions, the framework calls a new Claude session, prompting raw paragraph and related instructions to it, finally getting “Experiment Conclusions” in return. Breaking complex tasks into subtasks improves the performance of LLM-driven workflows Zhang and Gao (2023), which is the main reason for completing extraction and separation of conclusions in different Claude sessions.
Furthermore, we compared the extracted information by our approach, GPT-4, and Elicit (https://elicit.com/) over the same paper Bott et al. (2020). Elicit is an LLM fine-tuned on scholarly databases that can comprehend papers and extract related information Bail (2024), which could be a benchmark. As a result, whereas GPT-4 and Claude-3-Opus had a similar level of performance in this task, they both generated information of better quality than Elicit. As mentioned, Claude, rather than GPT, is chosen to extract information from the 319 papers to prevent potential data leakage.
4.2 Variant Generation
After the experiment settings and conclusions are extracted, the next step is to generate variants based on the true conclusion, since the goal is to see if the LLM could select the true one under distraction. Inspired by Luo et al. Luo et al. (2024)’s prediction of neuroscience results by LLMs, for each conclusion from an experiment, the framework prompts the original conclusion and its two variants to GPT: its reversed variant and unrelated variant. As a result, the framework will make the prediction by choosing one from three options.
Figure 5: Variant Generation Process
Specifically, as shown in Figure 5, the framework initially prompts the original conclusion to Claude, which follows the instructions to generate the reversed variant of the original conclusion. The reversed variant means that the direction of the conclusion is inverted. For example, if one conclusion is “receiving housing vouchers reduces quarterly employment rates,” its reversed variant is “receiving housing vouchers increases quarterly employment rates.” Next, the framework prompts both the original and reversed conclusions to Claude to generate the unrelated variant, which typically indicates that there is no correlation between entities of interest. Following the same example, the unrelated variant is “There is no relationship between receiving housing vouchers and quarterly employment rates.”
4.3 Prediction
In the final stage, our framework takes a field experiment’s conclusion and its two variants, and experiment settings as input and generates two parallel prediction prompts: Basic Prediction Prompt and Chain-of-Thought (CoT) Prediction Prompt, which are then prompted to GPT to get predictions by asking GPT to select one conclusion from three conclusions. Since CoT has proven to be capable of improving general performance of LLMs on downstream tasks Wei et al. (2022), it’s interesting to see how its effect on this prediction scenario.
Figure 6: Basic Prediction Prompt
Figure 7: Chain-of-Thought (CoT) Prediction Prompt
The basic prediction prompt is shown in Figure 6. It consists of a background information section, a question section, and necessary instructions. Specifically, the background information section contains the general goal of the experiment (such as exploring the impact of job training on income), treatments (such as receiving job training or not), experiment duration (such as seven weeks), outcomes (such as income), participant information (such as people seeking jobs in New England), and experiment workflow (such as when and how training was given and outcomes were recorded). All of these were extracted from a target paper, which is the same as bullet points “A” to “F” from Figure 4.
Following the background information, a question section is automatically generated and entered into the templates. Specifically, the original conclusion, its reversed variant, and its unrelated variant are shuffled and substitute the placeholders “option 1”, “option 2”, and “option 3”. Meanwhile, instructions tell GPT that the prediction of conclusions is under the context of the field experimental settings, asking GPT to choose one option from the three options as it deems correct.
As we are also interested in how CoT would improve such prediction, Figure 7 shows the CoT Prediction Prompt, which follows a similar logic as the basic prediction but integrates CoT strategies to boost the performance Wei et al. (2022). Initially, the framework prompts the experiment settings and three options of a conclusion to GPT, instructing it to think about decisive elements that help choose the correct option. Upon receiving the decisive elements from GPT, the framework prompts GPT to make the selection among three options to get a predicted conclusion. It is worth noting that the entire process is within the same GPT session for a conclusion.
Although either prompt strategy generates a prediction for a given input, LLMs are stochastic models, meaning that their responses may vary to the same prompt. To handle such randomness in experiment simulations, Leng & Yuan Leng and Yuan (2023) set the temperature to 0 and always get fixed responses from synthetic participants of lab experiments, which is a strategy to eliminate the stochasticity of LLMs totally. By contrast, Brand et al. Brand et al. (2023) repeated the same prompt 300 times and use the averaged number as the result of Willingness-to-Pay from customers role-played by LLMs. Here, we take the latter approach by incorporating the stochasticity of LLM outputs since this stochasticity of LLMs is similar to how the same human participant might respond differently when presented with the same instruction Demszky et al. (2023). Specifically, we repeat the same prediction prompt several times and calculate an average accuracy as the final result. For example, if the framework is running the basic prediction, a filled-out prompt based on Figure 6 will be repeated a given number of times to get a stable result. The determination of a proper repeat number will be further discussed in the Results section.
Parameter-wise, no fine-tuning is involved in any stage of the proposed framework, and all parameters of OpenAI API and Anthropic API are set to default. Whereas Horton Horton (2023) harnessed fine-tuned LLMs in lab experimental simulation to better follow instructions, Coda-Forno et al. Coda-Forno et al. (2024) simulated several human behaviors by LLMs without fine-tuning. The use of fine-tuned LLMs complicates the reproduction attempts since other researchers don’t have access to the same model in the existing papers Bail (2024). Additionally, avoiding fine-tuning LLMs saves computing resources and mitigates environmental impacts Ray (2023), especially when the pretraining of LLMs is enough to make them capable of downstream tasks Liu et al. (2023).
5 Results
In this section, we test the framework on 319 field experiments that contain a total of 1638 conclusions and discuss the performance.
“Conclusion Accuracy” and “Paper Accuracy” are used to evaluate the prediction performance under different settings. As shown in Figure 3, the framework generates a prompt for each conclusion. The generated prompt either follows the template in Figure 6 or 7, depending on which prompt strategy (basic or CoT) the framework applies. Each prompt instructs GPT to output a predicted conclusion. If the predicted conclusion matches the true conclusion, that attempt is counted as correct. As aforementioned, the framework repeats such an attempt for a set number of times for each conclusion to get a stable result. Therefore, we define “Conclusion Accuracy” as the percentage of correct predictions among a set number of attempts.
Conclusion Accuracy=Number of correct predictionsTotal number of $\text{Conclusion Accuracy}=\frac{\text{Number of correct predictions}}{\text{Total number of predictions}}\times 100\%$
Given that a field experiment may contain multiple conclusions, “Paper Accuracy” is the average of all “Conclusion Accuracy” within a paper.
Paper $\text{Paper Accuracy}=\frac{1}{N}\sum_{i=1}^{N}\text{Conclusion Accuracy}_{i}$
5.1 Prediction Performance Overview
GPT4-turbo Basic CoT
Conclusion Accuracy Paper Accuracy Conclusion Accuracy Paper Accuracy
Repeat = 10 62% 63% 73% 73%
Repeat = 20 61% 62% 73% 74%
Repeat = 30 61% 62% 73% 74%
Sample Size: 1638 conclusions from 319 papers
Table 1: Prediction Accuracy under Different Repeats
Repeat = 20 Basic CoT
Conclusion Accuracy Paper Accuracy Conclusion Accuracy Paper Accuracy
GPT3-turbo 60% 61% 66% 67%
GPT4-turbo 61% 62% 73% 74%
GPT4o 73% 74% 76% 77%
Sample Size: 1638 conclusions from 319 papers
Table 2: Prediction Accuracy under 20 Repeats by GPT Models
Table 1 reports two types of accuracy for both strategies under different repeats by GPT4-turbo (gpt-4-turbo-2024-04-09). The best results for the basic strategy are obtained under 10 repeats, which are 62% for conclusion accuracy and 63% for paper accuracy. By comparison, the best CoT results are 73% and 74% under either 20 or 30 repeats, which are generally 17% higher than basic results and aligns with prior literature on boosting LLM performance by CoT Wei et al. (2022). Meanwhile, another key observation is the results show a certain degree of invariability to repeat numbers. Specifically, the results across different numbers of attempts don’t seem to change under either the basic or CoT strategy. Given that more repeats result in higher time and monetary cost, we chose 20 repeats for further evaluation.
Fixing the attempts to 20, Table 2 reports results by different GPT models. GPT3-turbo, GPT4-turbo, and GPT4o here point to gpt-3.5-turbo-0125, gpt-4-turbo-2024-04-09, and gpt-4o-2024-11-20. They are the most representative models of their respective series, and their release dates are roughly six months apart from one another, which may reveal invariant features of LLMs in such a prediction task.
One key observation from Table 2 is that prediction performance steadily improves as LLMs iterate, though the rate of improvement decreases over time. According to Table 2, the best result under 20 repeats is achieved by GPT4o under CoT prompt strategy, which is a conclusion accuracy of 76% and a paper accuracy of 77%. Specifically, on average, GPT4o is able to predict a conclusion in 76% of the 20 repeated prediction attempts, and it also predicts 77% of the outcomes correctly for each paper. While the conclusion accuracy of GPT4o under CoT strategy is 15% higher than the CoT result of GPT3-turbo, it is only 4% higher than the CoT result of GPT4-turbo, indicating the improvement from model iteration becomes harder for this task.
Another interesting finding from Table 2 is the boosting effect of CoT on performance varies on models for the experiment conclusion prediction task. Specifically, the largest improvement is around 20% on GPT4-turbo, while the least improvement is around 4% on GPT4o. However, it’s unsafe to conclude that CoT boosting is weaker on newer models since CoT improves the accuracy by 10% on the oldest model tested, GPT3-turbo.
In summary, our findings indicate that incorporating a CoT strategy generally enhances predictive performance, whereas increasing the number of repetitions does not produce a significant effect. Furthermore, iterative refinements to LLMs consistently improve performance, although the rate of improvement diminishes over time. Finally, the performance gains attributable to CoT appear to be model-sensitive.
5.2 Skewness in Prediction Results
Figure 8: Conclusion Accuracy Distribution (GPT4o)
Figure 9: Paper Accuracy Distribution (GPT4o)
Two histograms (Figure 8 and 9) show the distribution of conclusion accuracy and paper accuracy obtained by GPT4o CoT prediction with 20 repeats, which is the best available result from Table 2.
As shown in Figure 8, the distribution of prediction accuracy exhibits a pronounced skewness, with a significant concentration of samples in the 90%-100% accuracy range (1165 conclusions) and another notable cluster in the 0%-10% accuracy range (278 conclusions). This indicates that the model’s performance is highly polarized, where certain conclusions are predicted with near-perfect accuracy, while others are almost wrongly predicted entirely. The similar pattern is also obvious in Figure 9 as the shape of distribution indicates that the framework performs well in a significant portion of papers.
Inspired by these unbalanced results, we closely examined the topics of each conclusion based on its experimental context. Leveraging LLMs’ ability in annotating Tan et al. (2024), we prompted Claude to label topic components of each conclusion under the context of the experiment. As a result, each conclusion got a binary vector containing only 0s and 1s like one-hot encoding vector, indicating the most related topic of it. Those topics include gender difference, socioeconomic status, ethnicity and social norms. Simiarly, Claude also decided the sentiment of each conclusion, either positive, negative, or neutral.
We find that the LLM-based prediction of experimental conclusions is sensitive to both topics and sentiments toward topics. For instance, when the conclusion is related to gender, and its context is favorable to males or detrimental to females, the conclusion accuracy is 74%. By contrast, when the context of the conclusion is favorable to females or detrimental to males, our model can make a obviously higher accuracy at over 87%. Similarly, the prediction accuracy is lower towards ethnicity-related conclusions with a negative sentiment at 80%, compared to the nearly perfect 94% of ethnicity-related conclusions with a positive sentiment. However, the LLM underperforms in the context of social norms (74%) and socioeconomic status (78%) regardless of the sentiment. In summary, we find LLMs have inclination or bias when predicting conclusions associated with specific elements because it does not fully capture mechanisms and merely make blind guesses based on their bias. This reveals scenarios on which the LLM-based field experiment simulation has limitations.
5.3 Robustness test with Manual Screening
Given that most steps in the proposed framework are automated, concerns naturally arise regarding the validity of these automated processes. To address these concerns, we conducted three manual screenings to verify the results. First, we examined whether the extracted experiment settings (Figure 3) inadvertently include genuine conclusions. Second, we checked whether the extracted conclusions align with those reported in the original papers. Finally, we assessed whether the generated variants of the conclusions (Figure 5) match our expectations.
Repeat = 20 Basic CoT
Conclusion Accuracy Paper Accuracy Conclusion Accuracy Paper Accuracy
GPT3-turbo 61% 61% 68% 67%
GPT4-turbo 66% 66% 76% 76%
GPT4o 75% 75% 78% 78%
Refined and Dequantified Sample Size: 1261 conclusions from 276 papers
Table 3: Prediction Accuracy of Refined and Dequantified samples under 20 Repeats by GPT Models
The first and third screenings revealed no issues. However, the second screening, which examined the alignment of extracted conclusions, identified 377 conclusions as either incomprehensible or nonexistent in the original texts. Consequently, after the three screenings, 1261 conclusions and 276 papers remained and were deemed valid for our purposes, reflecting approximately 77% of the initial sample size.
Additionally, 86 out of 1261 conclusions were dequantified, as existing studies suggest that predicting the magnitude of experimental outcomes remains challenging at this stage Manning et al. (2024). Given our focus on predicting the direction of experimental conclusions rather than the magnitude, conclusions specifying precise numerical treatment effects were reformulated in a dequantified manner. For instance, a conclusion such as “Job training increases income by 30%” was revised to “Job training increases income.”
Table 3 presents the aggregated findings of 1261 conclusions drawn from 276 papers. The overall performance improves marginally by approximately 2% percentage points relative to Table 2, with GPT4o retaining the highest accuracy at 78%. All observations discussed in Section 5.1 remain valid, and the skewness noted in Section 5.2 is likewise evident. Although manual screening reveals a few minor flaws in the proposed automated framework, the core conclusions—including prediction accuracy, polarization of results, model-sensitivity in CoT processes, the effect of model iteration on outcomes, and bias constraints—persist in manually refined samples, highlighting a robust reliability of the framework.
5.4 Handling Data Memorization
Figure 10: Paper Accuracy by Year
Data memorization is a common concern in simulating experiments with LLMs. If the results given by the LLM are from its memory of training data instead of reasoning, the proposed idea has no instructional value as pilot testing for field experiments Horton (2023). As revealed in its documents, the training data cutoff of gpt-4o-2024-11-20 is October 2023. Therefore, the report also reports accuracy by year and the standalone result for field experiments published in 2024. Notably, field experiments published after 2024 would be less likely to be included in the training data. Figure 10 shows that the performance on papers in 2024 is higher than any other, which we attribute to the small sample size of papers in 2024. Besides, for papers published from 2000 to 2023, paper accuracy of older field experiments is slightly better than more recent papers. Anyway, such a slow downward trend followed by a sudden rise in 2024 for paper accuracy is insufficient evidence of severe data memorization that denies GPT’s ability to predict field experimental conclusions.
6 Conclusions
In this paper, we proposed an LLM-powered framework that automatically extracts information from existing papers and predicts field experimental conclusions. To the best of our knowledge, this is the first paper to provide an automated framework for predicting such conclusions. Rather than merely introducing the framework, this is also the first work that examines its fidelity on a large scale of samples and clearly delineates infeasible scenarios in the experimental simulation literature. Building on these foundational efforts, our new results indicate that incorporating a CoT strategy generally enhances predictive performance, whereas merely increasing the number of repetitions does not produce a significant effect. Furthermore, iterative refinements to LLMs consistently improve performance, although the rate of improvement diminishes over time. Notably, the performance gains attributable to CoT appear to be model-sensitive. Taken together, these findings underscore the potential of the LLM-driven frameworks in advancing automated predictions for field experiments while also clarifying the practical constraints that guide their effective use.